OpenStudy (anonymous):

Solve the equation for x. x3 = 64

OpenStudy (anonymous):

calculate for the cube root of both sides

OpenStudy (anonymous):

\[\large \sqrt[3]{64}\]

OpenStudy (anonymous):

what is da answer

OpenStudy (anonymous):

\[\Large \sqrt[3]{64}=\sqrt[3]{2*32}=\sqrt[3]{2*2*16}=\sqrt[3]{2*2*2*8}=\sqrt[3]{2*2*2*2*4}=\sqrt[3]{2*2*2*2*2*2}=2*2=\]

OpenStudy (anonymous):

\(\sqrt[3]{2*2*2*2*4}=\sqrt[3]{2*2*2*2*2*2}=2*2=4\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

only for clarification: we are just here to guide you FOR YOU to get the right answer, we are not supposed to provide answers :)